OpenStudy (anonymous):

is there a 3-digit number for which\[\overline{abc}=b^2+4ac\]

OpenStudy (experimentx):

what's that symbol?

OpenStudy (anonymous):

that means abc is a number

hartnn (hartnn):

like abc=100*a+10b+c ?

OpenStudy (anonymous):

yeah

OpenStudy (experimentx):

100*a+10b+c = b^2+4ac <-- this seems like determinant of ax^2+bx-c = 0

OpenStudy (anonymous):

\[a,b,c \in \text{{0,1,...,9}}\]

OpenStudy (callisto):

a cannot be 0. Otherwise, it's not a 3-digit number.

OpenStudy (anonymous):

yeah a can not be 0

OpenStudy (anonymous):

RHS is limited

OpenStudy (callisto):

a, b, c are distinct, right?

OpenStudy (anonymous):

they can be same

OpenStudy (experimentx):

100*a+10b+ (c - b^2-4ac) = 0 (10 * 2a + b)^2 = b^2 - 4a(c - b^2-4ac)

OpenStudy (experimentx):

this is quadratic in a

OpenStudy (anonymous):

how u bring a^2 to work

OpenStudy (experimentx):

100*a+10b+ (c - b^2-4ac) = 0 x^2a+xb + (c - b^2-4ac) = 0, x=10

OpenStudy (experimentx):

perhaps we can use quadratic property and uniqueness to reduce things ... looks like this is too long and not my stuff. I should be doing PDE !!

OpenStudy (anonymous):

lol...man a is limited

OpenStudy (experimentx):

cannot have double value .. so discriminant should be zero.

OpenStudy (anonymous):

juantweaver started good

OpenStudy (dumbcow):

no number exists...the only possible chance is when a=1 but then there is no integer solution for b,c so both sides equal

OpenStudy (anonymous):

rewrite equation as \[100*a+c-4ac=b^2-10b\] then prove that LHS is always positive and RHS is always negative. impossible

OpenStudy (dumbcow):

i like that proof ^^

OpenStudy (anonymous):

that one is nice too juant

OpenStudy (anonymous):

only possibility for a is 1 -> only possibility for b -> 8 and 9 -> no integer value for c